# CSF: Fixed-outline Floorplanning Based on the Conjugate Subgradient Algorithm Assisted by Q-Learning

### v1

@@ IV-D2 Legalization based on the CSAQ
7:              Set $R(s_{j},a_{i})=0$ for all $i\in\{1,\dots,m\}$;
8:              $\mathbf{d}_{j,0}=\mathbf{0}$, $\mathbf{ind}^{0}_{j}=\mathbf{ind}_{j}$, $\mathbf{g}_{j,0}\in\partial f(\mathbf{ind}_{j}^{0})$;
@@
14:         Calculate the subgradient $\mathbf{g}_{j,k}\in\partial f(\mathbf{ind}_{j}^{k-1})$;
@@
19:         if $\mathbf{ind}_{j}^{k}$ is legal then
20:              succ=1, $\mathbf{FP^{*}}=\mathbf{ind}_{j}^{k}$;
21:              Break;
@@
24:     if succ=1 then
25:         Break;
26:     end if
27:end for
28:Update $\mathbf{FP}$ by the constraint graphs to get tight floorplans.
V Experimental Results
To verify the competitiveness of the proposed algorithm CSF, we perform numerical experiments on the GSRC and MCNC benchmarks. The I/O pads are fixed at the given coordinates for all of the test circuits, and all modules are set as hard modules. The algorithm is implemented in C++ and run on a 64-bit Windows laptop with 2.4 GHz Intel Core i7-13700H and 16-GB memory. The wirelength was evaluated by HPWL for all test cases. The aspect ratio $R$ is set as 1, unless otherwise specified, the $\gamma=15\%$ and the width and height are generated by
$W=\sqrt{(1+\gamma)\cdot A/R},\quad H=\sqrt{(1+\gamma)\cdot A\cdot R}.$ (14)
V-A The Positive Function of Q-learning
In the proposed CSAQ, the Q-learning strategy is introduced to automatically regulate the scale factor $c$ of the CSA. The adaptive regulation of $c$ contributes to better convergence of CSA in both the global floorplanning stage and the legalization stage.
V-A1 The performance of CSAQ-based global floorplanning
TABLE II: Performance comparison between CSA-based and CSAQ-based global floorplanning.
[…]
@@ V-A1 The performance of CSAQ-based global floorplanning
ami33 CSF-cc 0.19 0.41 0.72 73.3 133702 0.09 $\sim$
CSF-qc 0.81 0.53 1.61 93.3 131427
ami49 CSF-cc 0.13 0.08 0.35 96.6 916042 0.11 $\sim$
CSF-qc 0.41 0.10 0.65 100 924246
n100 CSF-cc 0.22 0.07 0.35 100 293202 $0.24$ $\sim$
CSF-qc 1.35 0.06 1.52 100 292880
n200 CSF-cc 0.63 0.13 0.82 100 524905 $0.03$ +
CSF-qc 4.27 0.18 4.62 100 523229
n300 CSF-cc 1.47 0.34 1.98 100 595735 $0.28$ $\sim$
CSF-qc 10.89 0.16 11.05 100 593784
To validate the performance improvement of CSAQ introduced by the Q-learning strategy, we compare two variants of CSF, 1) the CSF-cc implementing both global floorplanning and legalization by CSA and 2) the CSF-qc implementing global floorplanning and legalization by CSAQ and CSA, respectively. Due to the random characteristics of CSA, the performances are compared by averaged results of $30$ independent runs. The average runtime of a single run of global floorplanning ($t_{g}$), the average runtime of a single run of legalization ($t_{l}$), the average runtime of floorplanning ($t_{w}$), the success rate (SR) and the average wirelength (HPWL) are included in Table II. We further perform a one-side Wilcoxon rank-sum test for the optimization results of wirelength, and the p-values and the test results (R) with a significance level of $0.05$ are also demonstrated in Table II, where “+” indicates that CSF-qc significantly outperforms CSF-cc and $\sim$” means that the difference is not significant.
Compared with the CSF-cc, where CSA is not equipped with Q-learning, the CSF-qc also achieves an improved success rate on the MCNC benchmarks at the expense of a bit longer runtime. It also demonstrates that the better floorplanning results obtained by CSAQ can even shorten the runtime of legalization implemented by CSA. Although the Q-learning strategy is only introduced in the global floor planning stage, it is still effective in terms of the average wirelength of the final floorplan. However, the positive influence on the legalized floorplan is not very significant for the benchmarks except for the “n200” case.
V-A2 The performance of CSAQ-based legalization
The CSF-qc is further compared with the CSF-qq, where both global floorplanning and legalization are implemented by addressing (3) via CSAQ, and the results are included in Table III.
TABLE III: Performance comparison between CSA-based and QCSA-based legalization.
[…]
@@ V-A2 The performance of CSAQ-based legalization
ami33 CSF-qc 0.81 0.53 1.61 93.3 131427 $4.79\times 10^{-11}$ +
CSF-qq 0.75 0.98 1.88 100 84427
ami49 CSF-qc 0.41 0.10 0.65 100 924246 0.21 ~
CSF-qq 0.45 0.26 0.78 100 902566
n100 CSF-qc 1.35 0.06 1.52 100 292880 $7.92\times 10^{-5}$ +
CSF-qq 1.33 0.25 1.68 100 290207
n200 CSF-qc 4.27 0.18 4.62 100 523229 $9.30\times 10^{-7}$ +
CSF-qq 4.28 0.64 5.25 100 519860
n300 CSF-qc 10.89 0.16 11.05 100 593784 $4.76\times 10^{-4}$ +
CSF-qq 11.21 0.92 12.13 100 588951
@@
The results show that by employing the Q-learning strategy, the CSAQ leads to significantly better legalization results of the investigated benchmarks. The statistical test indicates that CSF-qq performs significantly better than CSF-qc on 4 of 5 problems. Although the test does not demonstrate significant superiority of CSF-qq in the ami49 case, it still beats CSF-qc in terms of the average wirelength.
V-B Performance comparison of legalization algorithms
Li et al. [12] proposed a CG-based legalization algorithm (LA) for large-scale fixed-outline floorplanning, and an improved variant (ILA) is proposed by us in Algorithm 6 to further refine the results of wirelength optimization. Motivated by the idea of legalizing the floorplan by addressing a nonsmooth optimization problem [22], we also proposed a CSAQ-based legalization algorithm (CLA) presented by Algorithm 7. To compare their performance, we generate the initial floorplan to be legalized by the CSAQ-based global floorplanning algorithm presented in Algorithm 4.
Note that the ILA involves a parameter $k$ that determines the number of candidate correction relationships to be selected probabilistically. In our experiments, we get two variants of ILA by setting $k=3$ with two probability distributions.
  • ILAm: $Pw_{1}=0.9$, $Pw_{2}=0.05$, $Pw_{3}=0.05$;
  • ILAs: $Pw_{1}=1$, $Pw_{2}=0$, $Pw_{3}=0$.
@@ V-B Performance comparison of legalization algorithms
For each benchmark problem, the white space is set to 20%, and 30 independent experiments are performed to get the statistical results presented in TABLE IV, where MinW and HWSD represent the minimum value and the standard deviation of the results of 50 independent runs, respectively.
TABLE IV: Comparison of Different legalization Methods
[…]
ami33 LA [12] 116384 89851 0.51 90 14852.88
ILAm 109153 86236 0.17 100 10306.27
ILAs 101496 86147 0.63 100 11993.17
CLA 83452 75783 0.08 100 5383.99
ami49 LA [12] 1563721 980543 0.33 96 253380.55
ILAm 1304216 921645 0.29 96 252591.71
ILAs 1292577 894563 1.04 90 181419.36
CLA 917899 806937 0.15 100 62604.82
n100 LA [12] 335610 310834 3.64 100 18157.76
ILAm 314903 286532 2.42 100 14945.68
ILAs 305016 286748 2.84 100 12775.31
CLA 290495 286277 1.42 100 2273.33
n200 LA [12] 636552 538590 22.22 100 29538.32
ILAm 568583 513892 12.17 100 22937.34
ILAs 560880 515428 12.67 100 25125.83
CLA 520160 516038 4.61 100 2391.86
n300 LA [12] 806155 588942 66.30 100 53314.69
ILAm 667439 581653 42.37 100 22937.34
ILAs 636666 583300 38.87 100 30192.93
CLA 591151 581712 11.29 100 4667.06
@@
The results show that the improved CG-based legalization algorithms (ILAm and ILAs) generally outperform the LA in terms of both the average wirelength and the minimum wirelength. Since ILAs always selects the candidate correction relationship with the smallest wirelength, it optimizes the wirelength better than ILAm. However, the heuristic selection strategy of ILAs sometimes leads to unpromising local convergence of the legalization process, resulting in a lower success rate for the ami49 circuit. Compared with the CG-based legalization methods including LA, ILAm and ILAs, the CLA can get legalized floorplans with the shorter wirelength, demonstrating its superiority in optimization of wirelength.
@@
From the perspectives of the average time and the success rate, ILAm has outstanding performance in terms of the average runtime, which validates that the probabilistic selection strategy for the removal of critical relationships contributes to a fast success of legalization. Meanwhile, the CLA gets legalized floorplans with the shortest runtime, and the significantly smaller HWSD values demonstrate its superior robustness regarding the wirelength of floorplan.
In conclusion, the improved CG-based algorithms (ILAs and ILAm) are capable of generating better floorplans compared to the LA, while the CSAQ-based algorithm can get fast and robust legal floorplans.
V-C Comparison between CSF-qq and other algorithms.
@@
The comparison in Section V-A demonstrates that the Q-learning strategy significantly improves the performance of the CSA-based global floorplanning and legalization, and the competitiveness of the CSAQ-based legalization is also validated in Section V-B. To further verify the performance of the CSA-based fixed-ouline floorplanning algorithm assisted by Q-learning, we compare the CSF-qq with some selected floorplanning algorithms, the Parquet-4.5 [3], the Fast-FA [5], the FFA-CD [22] and the Per-RMAP [17]. Among them, Parquet-4.5 and Fast-FA are based on the B*-tree coding structure of floorplan, and the optimization of wirelength is addressed by the simulated annealing algorithm; FFA-CD utilizes the distribution evolution algorithm based on the probability model to optimize the orientation of modules and employs CSA to optimize the corresponding coordinates of each module to obtain the final solution; Per-RMAP models the floorplanning problem as a feasibility-seeking problem(FSP), and then deals with the floorplanning problem by introducing a perturbed resettable strategy into the method of alternating projection(MAP). The experiment considers the fixed outline floorplanning of hard modules with a white space setting of $\gamma=15\%$, and the statistical results are shown in Table V. Note that the experimental data of Fast-FA and Per-RMAP are taken from [17], where the success rates are not presented for the benchmark circuits.
TABLE V: Performance Comparison of Selected Fixed-outline Floorplanning Algoirithms
[…]
ami33 Parquet-4.5 78650 0.16 60
Per-RMAP 63079 0.14 /
CSF-qq 83665 1.76 100
ami49 Parquet-4.5 962654 0.37 60
Per-RMAP 689296 1.70 /
CSF-qq 905653 0.75 100
n100 Parquet-4.5 334719 1.73 100
Fast-SA 287646 10.72 /
FFA-CD 293578 0.70 100
Per-RMAP 282596 1.01 /
CSF-qq 290156 1.78 100
n200 Parquet-4.5 620097 7.78 100
Fast-SA 516057 69.86 /
FFA-CD 521140 1.95 100
Per-RMAP 518722 2.93 /
CSF-qq 519211 5.81 100
n300 Parquet-4.5 768747 16.86 100
Fast-SA 603811 133.63 /
FFA-CD 588118 3.44 100
Per-RMAP 626061 4.11 /
CSF-qq 588866 13.10 100
The numerical results show that, compared with Parquet-4.5, the performance of CSF-qq in optimizing wirelength on GSRC has been improved by 13.3%~23.2%. In addition, CSF-qq far outperforms Fast-SA in terms of running time, and it is also superior in optimizing the wirelength for n300. When compared with the current advanced algorithms FFA-CD and Per-RMAP, CSF-qq remains competitive in both wirelength and running time on GSRC. Finally, we show the fixed-outline floorplanning results of a single independent experiment of CSF-qq on the GSRC benchmarks in Figure 1.

### v2

@@ 4.4.2 The legalization algorithm based on CSAQ
7:     Set $R(s_{j},a_{i})=0$ for all $i\in\{1,\dots,m\}$, initialize $c_{j}$;
8:     $\mathbf{d}_{j,0}=\mathbf{0}$, $\mathbf{Ind}^{0}_{j}=\mathbf{Ind}_{j}$, $\mathbf{g}_{j,0}\in\partial f_{l}(\mathbf{Ind}_{j}^{0})$;
@@
14:   Calculate the subgradient $\mathbf{g}_{j,k}\in\partial f_{l}(\mathbf{Ind}_{j}^{k-1})$;
@@
19:   if $f_{l}(\mathbf{Ind}_{j}^{k})=0$ then
20:     $\mathbf{Ind}^{*}=\mathbf{Ind}_{j}^{k}$;
21:     $\mathbf{return}$;
@@
24:end for
25:Update $\mathbf{Ind}^{*}$ by the constraint graphs to get tight floorplans.
5 Experimental Results
To verify the competitiveness of CSF, we perform numerical experiments on the GSRC and MCNC benchmarks, where the I/O pads are fixed at the given coordinates for all of the test circuits, and all modules are set as hard modules. The characteristics of the benchmarks are presented in Tab. 2, and the compared algorithms are implemented in C++ and run on a 64-bit Windows laptop with 2.4 GHz Intel Core i7-13700H and 16-GB memory.
Table 2: Characteristics of the Benchmarks.
Benchmark Instance # Modules # IO Pins # Pins # Nets Die Size ($\mu m^{2}$)
$s1$ $s2$
MCNC ami33 $33$ $42$ $480$ $123$ $1153.22\times 1153.22$ $2058\times 1463$
ami49 $49$ $22$ $931$ $408$ $6384.53\times 6384.53$ $7672\times 7840$
GSRC n100 $100$ $334$ $1873$ $885$ $454.341\times 454.341$ $800\times 800$
n200 $200$ $564$ $3599$ $1585$ $449.5\times 449.5$ $800\times 800$
n300 $300$ $569$ $4358$ $1893$ $560.487\times 560.487$ $800\times 800$
As presented in Tab. 2, the fixed-outline floorplanning is performed within two different settings of die size. For setting $s1$, the die sizes are generated by (6) with $R=1$ and $\gamma=15\%$. The size setting $s2$ is cited from Ref. [17] to validate the wirelength minimization capability of the CSF. By numerical experiments, we first show how the incorporated Q-learning enhances the performance of CSF, and then, compare it with several selected floorplanning algorithms. The comparison is performed for the cases without soft modules to highlight how the CSF efficiently addresses the great challenge of minimizing the wirelength within a fixed outline for the scenarios containing only hard modules.
In the proposed CSAQ, the Q-learning strategy is introduced to automatically regulate the scale factor $c$ of the CSA. Influenced by module sizes and the primary objectives of different stages, we design distinct action groups for updating the $c$ across various benchmarks. Specifically:
  • 1.
    For the global floorplanning stage:
    • (a)
      On GSRC benchmarks, the action groups are [8, 12, 15, 20, 25];
    • (b)
      For ami33, the action groups are [80, 100, 120, 140, 160];
    • (c)
      For ami49, the action groups are [130, 180, 220, 270, 330].
  • 2.
    For the legalization stage:
    • (a)
      On GSRC benchmarks, the action groups are [0.1, 0.8, 5, 10, 20];
    • (b)
      For ami33, the action groups are [1, 8, 30, 60, 90];
    • (c)
      For ami49, the action groups are [10, 50, 100, 150, 200].
5.1 The Positive Effects of Q-learning
The positive effects of Q-learning are validated by comparing the performance of different global floorplanning (GP) and legalization (LA) strategies, where the $\alpha_{0},\gamma_{0},m_{0}$ of (10) and (11) are 0.4, 0.8, 0.6 respectively, and the weights $\alpha,\lambda,\mu$ of (8) and (9) as well as the parameters $k_{t},k_{max},c_{0}$ of CSA and CSAQ are presented in Tab. 3. The fixed-outline die sizes are set as the $s1$ of Tab. 2.
Table 3: Parameter settings of CSA and CSAQ in different stage.
Stage Optimizer Benchmark $k_{t}$ $k_{max}$ $\alpha$ $\lambda$ $\mu$ $c_{0}$ $v$
GP CSA MCNC - 50 1 20 10 1000 10
GSRC - 50 1 20 100 100 100
CSAQ MCNC 10 50 1 20 10 330 10
GSRC 40 200 1 20 100 25 100
LA CSA MCNC - 2000 - 1 10 500 -
GSRC - 2000 - 1 10 50 -
CSAQ MCNC 100 2000 - 1 10 100 -
GSRC 100 5000 - 1 10 10 -
5.1.1 The performance of CSAQ-based global floorplanning
Table 4: Performance comparison between CSA- and CSAQ-based global floorplanning.
Circuit Method Runtime SR($\%$) HPWL($\mu m$) IR($\%$) p-value R
@@ 5.1.1 The performance of CSAQ-based global floorplanning
ami33 CSF-$c$ 0.008 0.006 0.23 73.3 133702 1.70 0.09 $\sim$
CSF-qc 0.031 0.005 0.66 80 131427
ami49 CSF-$c$ 0.019 0.018 0.31 96.6 916042 -0.89 0.11 $\sim$
CSF-qc 0.11 0.016 0.45 100 924246
n100 CSF-$c$ 0.22 0.07 0.35 83.3 293202 0.14 $0.24$ $\sim$
CSF-qc 1.35 0.06 1.52 100 292780
n200 CSF-$c$ 0.63 0.16 0.82 96.6 524905 0.32 $0.03$ +
CSF-qc 4.27 0.13 4.62 100 523229
n300 CSF-$c$ 1.47 0.34 1.98 90 595735 0.33 $0.28$ $\sim$
CSF-qc 10.89 0.21 11.05 100 593784
To validate the promising effect of CSAQ on the global floorplanning stage, we compare two variants of CSF, 1) the CSF-cc implementing both global floorplanning and legalization by CSA and 2) the CSF-qc implementing global floorplanning and legalization by CSAQ and CSA, respectively. Due to the random characteristics of CSA, the performances are compared by averaged results of $30$ independent runs. The average runtime of a single run of global floorplanning ($t_{g}$), the average runtime of a single run of legalization ($t_{l}$), the average runtime to complete the entire floorplanning ($t_{w}$), the success rate (SR), the average wirelength (HPWL) and the improvement ratio(IR) of the HPWL are included in Table 4. Moreover, the one-sided Wilcoxon rank-sum test is performed based on the wirelength data of 30 independent runs, and the p-values and the test results (R) with a significance level of $0.05$ are also demonstrated in Table 4, where “+” indicates that CSF-qc significantly outperforms CSF-cc and $\sim$” means that the difference is not significant.
Compared to CSF-cc, CSF-qc demonstrates an enhanced success rate at the cost of a marginally longer $t_{g}$. However, the runtime of leglaization $t_{l}$ is shorter, which indicates that the better floorplanning results obtained by CSAQ can even shorten the runtime of legalization implemented by CSA. Consequently, the CSF-qc is competitive to CSF-cc in the terms of averaged wirelength for most of the test benchmarks. However, the CSAQ-assisted GP alone cannot improve the floorplanning results to a large extent, and the hypothesis test shows that CSF-qc only outperforms CSF-cc significantly on the n200 case.
5.1.2 The superiority of CSAQ-based legalization to CSA-based legalization
The CSF-qc is further compared with the CSF-qq, where both global floorplanning and legalization are implemented by CSAQ. The results are presented in Table 5.
Table 5: Performance comparison between CSA- and QCSA-based legalization.
Circuit Method Runtime SR($\%$) HPWL($\mu m$) IR($\%$) p-value R
@@ 5.1.2 The superiority of CSAQ-based legalization to CSA-based legalization
ami33 CSF-qc 0.031 0.007 0.66 80 131427 36.52 $4.79\times 10^{-11}$ +
CSF-$q$ 0.032 0.11 0.91 83.3 83427
ami49 CSF-qc 0.11 0.017 0.45 100 924246 2.02 0.21 ˜
CSF-$q$ 0.10 0.13 0.58 100 905566
n100 CSF-qc 1.35 0.06 1.52 100 292880 0.93 $7.92\times 10^{-5}$ +
CSF-$q$ 1.33 0.25 1.76 100 290156
n200 CSF-qc 4.27 0.18 4.62 100 523229 0.77 $9.30\times 10^{-7}$ +
CSF-$q$ 4.28 0.64 5.35 100 519211
n300 CSF-qc 10.89 0.16 11.05 100 593784 0.81 $4.76\times 10^{-4}$ +
CSF-$q$ 11.21 0.92 12.40 100 588951
@@
It is demonstrated that by employing the Q-learning strategy, the CSAQ leads to significantly better legalization results at the expense of slightly increased runtime. The statistical test indicates that CSF-qq outperforms CSF-qc on four benchmark circuits. Although the test does not demonstrate significant superiority of CSF-qq in the ami49 case, it still beats CSF-qc in terms of the average wirelength.
5.2 Comparison between CG- and QCSA-based legalization algorithms
To further confirm the superiority of the CSAQ-based legalization algorithm, we compare the LA-CSAQ (Algorithm 7) with the LA-CG  [12] by starting the legalization with initial floorplans generated by the CSAQ based global floorplanning algorithm $GFloorplan(f_{g},\mathbf{Pop})$(Algorithm 4). Meanwhile, two variants of the ILA-CG (Algorithm 6) are investigated to highlight the superiority of LA-CSAQ. Both variants of the ILA-CG set $k=3$ with two different probability distributions.
  • 1.
    ILA-CGm: $Pw_{1}=0.9$, $Pw_{2}=0.05$, $Pw_{3}=0.05$;
  • 2.
    ILA-CGs: $Pw_{1}=1$, $Pw_{2}=0$, $Pw_{3}=0$.
@@ 5.2 Comparison between CG- and QCSA-based legalization algorithms
The statistical results of $30$ independent runs are included in Table 6, where MinW and HWSD represent the wirelength of minimum value and the standard deviation, respectively.
Table 6: Performance Comparison Between CG and QCSA-based Legalization Algorithms
Circuit Method HPWL($\mu m$) MinW($\mu m$) CPU($s$) SR($\%$) HWSD($\mu m$)
ami33 LA-CG [12] 119625 88457 2.35 43.3 10342.88
ILA-CGm 109235 81047 1.26 80 104487.27
ILA-CGs 103136 79173 2.45 30 16993.17
LA-CSAQ 83427 75783 0.91 83.3 5383.99
ami49 LA-CG [12] 1626100 1275200 1.03 100 147349.55
ILA-CGm 1481960 1064360 1.23 100 164380.71
ILA-CGs 1422600 928432 2.73 100 194858.36
LA-CSAQ 905566 806937 0.58 100 62604.82
n100 LA-CG [12] 389435 369092 37.58 60 18157.76
ILA-CGm 363257 323760 31.87 53.3 14945.68
ILA-CGs 359603 312565 35.37 60 15775.31
LA-CSAQ 290156 286277 1.76 100 2273.33
n200 LA-CG [12] 718581 698852 187.32 10 24538.32
ILA-CGm 708955 682711 163.25 30 21937.34
ILA-CGs 699861 666166 12.67 50 19825.83
LA-CSAQ 519211 516038 5.35 100 2391.86
n300 LA-CG [12] - - - 0 -
ILA-CGm - - - 0 -
ILA-CGs - - - 0 -
LA-CSAQ 588951 580712 12.40 100 4667.06
@@
The results show that the improved CG-based legalization algorithms (ILA-CGm and ILA-CGs) generally outperform the LA-CG in terms of both the average wirelength and the minimum wirelength. Since ILA-CGs always selects the candidate correction relationship contributing to the smallest wirelength, it optimizes the wirelength better than ILA-CGm. Compared with the CG-based legalization methods including LA-CG, ILA-CGm and ILA-CGs, the LA-CSAQ can get legal floorplans with the shorter wirelength, demonstrating its superiority in optimization of wirelength.
@@
From the perspectives of the average time and the success rate, the LA-CSAQ gets legal floorplans with the shortest runtime and the highest success rate, and the significantly smaller HWSD values demonstrate its superior robustness regarding the wirelength of floorplan. Although both ILA-CGm and ILA-CGs outperform LA-CG on circuits ami33, ami49, n100 and n200, none of the CG-based legalization algorithms can eliminate the constraint violations for the n300 circuit, which further confirms the superior performance of LA-CSAQ in the task of floorplan legalization.
5.3 Comparison between CSF and the state-of-the-art floorplanning algorithms
@@
The competitiveness of the proposed CSAQ-based floorplanning algorithm is further validated by comparing the CSF-qq (where both global floorplanning and legalization are addressed by CSAQ) with some selected floorplanning algorithms, the Parquet-4.5 [3], the Fast-FA [5], the FFA-CD [22] and the Per-RMAP [17]. Among them, Parquet-4.5 and Fast-FA are based on the B*-tree coding structure of floorplan, and the optimization of wirelength is addressed by the simulated annealing algorithm; FFA-CD utilizes the distribution evolution algorithm based on the probability model to optimize the orientation of modules and employs CSA to optimize the module coordinates; Per-RMAP models the floorplanning problem as a feasibility-seeking problem, and then deals with the floorplanning problem by introducing a perturbed resettable strategy into the method of alternating projection.
For the fixed-outline floorplanning scenario, the introduction of soft module is beneficial to achieve legal floorplans with short wirelength. However, this paper is dedicated to the scenario of fixed-outline floorplanning without soft modules. Note that the experimental data of Fast-FA and Per-RMAP are taken from Ref. [17], in which the success rates of the benchmark circuits are not presented.
5.3.1 Floorplanning within compact layout regions
We first validate the performance of CSF-qq on floorplanning scenarios with a tightly arranged layout region by comparing it with Parquet-4.5 and FFA-CD. The layout region is set as setting $s1$ presented in Tab. 2. The generated floorplans of CSF-qq on the GSRC benchmarks are illustrated in Fig. 2, and the statistical results of 30 independent runs are presented in Table 7, where IR represents the wirelength improvement ratio of CSF-qq compared to other methods.